OpenStudy (anonymous):

Choose the missing step in the given solution to the inequality −6x − 10 > 14 + 2x. −6x − 10 > 14 + 2x _______________ −8x > 24 x < −3 −4x − 10 > 14 −8x − 10 > 14 −4x − 10 < 14 −8x − 10 < 14

OpenStudy (xhawaiiangirlx):

I believe it is step 1. Because : |2x - 4| - 3 = 3 |2x - 4| = 3 + 3 (not 3 - 3)

OpenStudy (anonymous):

thank you sooo much can you please help me with some others?

OpenStudy (anonymous):

@XhawaiiangirlX

OpenStudy (texaschic101):

missing step : -8x - 10 > 14

OpenStudy (anonymous):

Solve for L: A = L ⋅ W ⋅ H L = A times W over H L = W times H over A L = A over quantity W times H L = W over quantity H times A

OpenStudy (texaschic101):

divide both sides by WH and what do you get ?

OpenStudy (anonymous):

a*w*h=l

OpenStudy (texaschic101):

A = LWH -- divide by WH A/WH = L

OpenStudy (texaschic101):

third one is your answer. Understand ?

OpenStudy (anonymous):

yes thank you sooo much!!!! can you please help me with a few more

OpenStudy (texaschic101):

I can try

OpenStudy (texaschic101):

ask away

OpenStudy (anonymous):

The table below shows two equations: Equation 1 |3x − 1| + 7 = 2 Equation 2 |2x + 1| + 4 = 3 Which statement is true about the solution to the two equations? Equation 1 and equation 2 have no solutions. Equation 1 has no solution and equation 2 has solutions x = 0, 1. The solutions to equation 1 are x = −1.3, 2 and equation 2 has no solution. The solutions to equation 1 are x = −1.3, 2 and equation 2 has solutions x = 0, 1.

OpenStudy (anonymous):

?

OpenStudy (texaschic101):

| 3x - 1 | + 7 = 2 --- subtract 7 from both sides | 3x - 1 | = 2 - 7 | 3x - 1 | = -5 3x - 1 = -5 -(3x - 1) = -5 3x = -5 + 1 -3x + 1 = -5 3x = -4 -3x = -5 - 1 x = -4/3 -3x = -6 x = 2 so x = -4/3(or -1.3) and 2 | 2x + 1| + 4 = 3 | 2x + 1| = 3 - 4 | 2x + 1 | = -1 2x + 1 = -1 -(2x + 1) = -1 2x = -1-1 -2x - 1 = -1 2x = -2 -2x = -1 + 1 x = -1 -2x = 0 x = 0 so we have x = -1 and 0 In that last answer choice...is it supposed to be (in eq 2)..x = 0 and -1 ?

OpenStudy (texaschic101):

I am not getting any of your answer choices...closest I am coming to is the last one...but I am getting -1 instead of 1

OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality x − 3 < −11 or x + 5 greater than or equal to 14. A number line with a closed circle on −8, shading to the left, and an open circle on 9, shading to the right A number line with an open circle on −8, a closed circle on 9, and shading in between A number line with an open circle on −8, shading to the left, and a closed circle on 9, shading to the right A number line with a closed circle on −8, an open circle on 9, and shading in between

OpenStudy (texaschic101):

x - 3 < -11 x < -11 + 3 x < - 8 x + 5 >= 14 x >= 14 - 5 x >= 9 so we have a number line with an open circle on -8 shading to the left and a closed circle on 9, shaded to the right

OpenStudy (anonymous):

In a certain county, the number of charter schools is 6 less than twice the number of alternative schools. We know that there are 52 charter schools in the county. How many alternative schools are in the county? 58 44 29 23

OpenStudy (texaschic101):

c = 2a - 6 c = 52 so we sub 52 = 2a - 6 52 + 6 = 2a 58 = 2a 58/2 = a 29 = a....29 alternative schools

OpenStudy (anonymous):

the last one Solve for x. −ax + 3b > 5 x > the quantity 3 times b minus 5 all over a x > the quantity 5 minus 3 times b all over negative a x < the quantity 3 times b plus 5 all over a x < the quantity negative 3 times b plus 5 all over negative a

OpenStudy (texaschic101):

-ax + 3b > 5...solving for x -ax > -3b + 5 x < (-3b+5)/-a....last answer choice

OpenStudy (anonymous):

thank you soooo much, you have no clue how much you helped me, you are amazing!!!!

OpenStudy (texaschic101):

lol...I try...happy to help :)

OpenStudy (anonymous):

Solve for H: A = L ⋅ W ⋅ H H = A times W over L H = W times L over A H = L times A over W H = A over quantity L times W

OpenStudy (anonymous):

@texaschic101